OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, -7) and (6, -5)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (sheraz12345):

Apply formula y2-y1/x2-x1

OpenStudy (anonymous):

A.1/6 B.- 1 C.1/2 D.undefined

hartnn (hartnn):

x1 = -6 y1 =-7 x2 =6 y2 =-5 what is y2 - y1 = ...? x2-x1 =... ?

OpenStudy (anonymous):

the answer is A

hartnn (hartnn):

you are correct :)